A bus moves from rest with a uniform acceleration of 2 m/s2 for the first 10 s. It then accelerate at a uniform rate of 1 m/s2 for another 15 s. It continues at constant speed for 70 s and finally comes to rest in 20 s by uniform deceleration. Draw its velocity time graph and calculate total distance covered.


1
Expert's answer
2020-10-27T08:24:04-0400

Solution

Graph can be drawn as



Are of v-t curve will give displacement

So

Total are of OABCD IS

"S=\frac{1}{2}\times20\times10+((20\times15)+\frac{1}{2}\times15\times15) \\+35\times70+\frac{1}{2}\times20\times35"

"S=3312.5m"

So distance travelled by bus is 3312.5m